# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ Abstract
The moduli space of stable Higgs bundles of degree 00 is equipped with the hyperkähler metric, called the Hitchin metric. On the locus where the spectral curves are smooth, there is the hyperkähler metric called the semi-flat metric, associated with the algebraic integrable systems with the Hitchin section. We prove the exponentially rapid decay of the difference between the Hitchin metric and the semi-flat metric along the ray $(E,t\theta)$ as $t\to\infty$.
@@ 1.1.1 Hitchin metric
Let $X$ be a compact Riemann surface with a conformal metric $g_{X}$. In this paper, a Higgs bundle on $X$ is a holomorphic vector bundle $E$ with a holomorphic section $\theta$ of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$, where $K_{X}$ denotes the canonical bundle of $X$. The study of the moduli space of Higgs bundles was initiated by Hitchin in [14]. Let $\mathcal{M}_{H}$ denote the moduli space of stable Higgs bundles of degree 00 and rank $n$ on $X$. It is naturally equipped with the integrable almost complex structure $I$. Moreover, $\mathcal{M}_{H}$ is equipped with the naturally defined hyperkähler structure, i.e., the Riemannian metric $g_{H,{\mathbb{R}}}$ and the three integrable almost complex structures $I$, $J$ and $K$ such that (i) the associated $2$-forms $\omega_{T}(\cdot,\cdot)=g_{H,{\mathbb{R}}}(\cdot,T\cdot)$ $(T=I,J,K)$ are closed, (ii) $IJK=-1$. Let $g_{H}$ denote the associated Kähler metric of $(\mathcal{M}_{H},I)$, i.e., $g_{H}=g_{H,{\mathbb{R}}}+\sqrt{-1}\omega_{I}$. The Riemannian metric $g_{H,{\mathbb{R}}}$ and the Kähler metric $g_{H}$ are called the Hitchin metric.
By the theorem of Hitchin and Simpson [14, 25], because $(E,\theta)\in\mathcal{M}_{H}$ is stable of degree 00, it carries a harmonic metric $h$, i.e., $h$ is a Hermitian metric of $E$ satisfying
@@
Here, $R(h)$ denotes the curvature of the Chern connection of $(E,h)$, and $\theta^{\dagger}_{h}$ denotes the adjoint of $\theta$ with respect to $h$. The induced Hermitian metric on $\mathop{\rm End}\nolimits(E)$ is also denoted by $h$. The tangent space $T_{(E,\theta)}\mathcal{M}$ is identified with the space of harmonic $1$-forms of the harmonic bundle $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$. The Hermitian metric $(g_{H})_{|(E,\theta)}$ is identified with the $L^{2}$-metric on the space of harmonic $1$-forms. (For example, see §4.2.3.) Namely, for harmonic $1$-forms $\tau$ of $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$, let $[\tau]\in T_{(E,\theta)}\mathcal{M}_{H}$ denote the corresponding element, and then we have
@@ 1.1.2 Semi-flat metric
We set $\mathcal{A}_{H}=\bigoplus_{j=1}^{n}H^{0}(X,K_{X}^{j})$. Let $\pi_{0}:T^{\ast}X\to X$ denote the projection. Let $\eta$ denote the tautological section of $\pi_{0}^{\ast}(K_{X})$. For any $a=(a_{j})\in\mathcal{A}_{H}$, we obtain the section $\xi_{a}=\eta^{n}+\sum_{j=1}^{n}(-1)a_{j}\eta^{n-j}$ of $\pi_{0}^{\ast}(K_{X}^{\otimes n})$. Let $\Sigma_{a}$ denote the 00-scheme of the section $\xi_{a}$. Such a curve is called a spectral curve. Let $\mathcal{A}_{H}^{\prime}\subset\mathcal{A}_{H}$ denote the Zariski open subset of $a\in\mathcal{A}_{H}$ such that $\Sigma_{a}$ are smooth. It is known that $\mathcal{A}_{H}^{\prime}$ is non-empty.
@@
It is independent of the choice of a Hitchin section. Because the fiber $\Phi_{H}^{-1}(\Phi_{H}(E,\theta))$ is identified with a connected component of the Picard variety of $\Sigma_{\theta}$, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt v}}$ is naturally identified with $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$. The moduli space $\mathcal{M}_{H}^{\prime}$ is equipped with a complex symplectic structure. Because $\Phi_{H}^{-1}(\Phi_{H}(E,\theta))$ is Lagrangian, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt v}}$ and $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt h}}$ are mutually dual. Hence, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt h}}$ is naturally identified with $H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$. The space $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\oplus H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ is identified with the space of harmonic $1$-forms on $\Sigma_{\theta}$. The Hermitian metric $(g_{\mathop{\rm sf}\nolimits})_{(E,\theta)}$ is identified with the $L^{2}$-metric on the space of harmonic $1$-forms on $\Sigma_{\theta}$. (See Proposition 4.6.) Namely, for harmonic $1$-forms $\tau$ on $\Sigma_{\theta}$, let $[\tau]\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\oplus H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ denote the corresponding element, and then we have
@@ 1.1.3 A weak version of a conjecture of Gaiotto-Moore-Neitzke
A previous work was done by Fredrickson in [5] on the basis of her another work [6], the works of Mazzeo-Swoboda-Weiss-Witt [18, 19] and interesting ideas of Dumas-Neitzke [4]. More precisely, for each $(E,\theta)\in\mathcal{M}_{H}^{\prime}$, let $\mathop{\rm Crit}\nolimits(\theta)\subset\Sigma_{\theta}$ denote the set of the critical points of the projection $\pi:\Sigma_{\theta}\to X$. For each $Q\in\mathop{\rm Crit}\nolimits(\theta)$, let $r_{Q}$ denote the ramification index of $\pi$ at $Q$. We obtain the Zariski open subset $\mathcal{M}_{H}^{\prime\prime}$ of $(E,\theta)\in\mathcal{M}_{H}^{\prime}$ such that $r_{Q}\leq 2$ for any $Q\in\mathop{\rm Crit}\nolimits(\theta)$. We have $\mathcal{M}^{\prime}_{H}=\mathcal{M}_{H}^{\prime\prime}$ if $n=2$ but $\mathcal{M}_{H}^{\prime}\neq\mathcal{M}_{H}^{\prime\prime}$ if $n\geq 3$, in general. The conjectural estimate (1) has been studied for any $(E,\theta)\in\mathcal{M}_{H}^{\prime\prime}$ in [5]. We should also note that a similar estimate has been studied by Fredrickson-Mazzeo-Swoboda-Weiss [7] in the case of parabolic Higgs bundles of rank $2$. Chen-Li [2] studied the rank $2$ but more singular case by using the methods in [5, 7]. In [15], Holdt studied the original approach of [9, 10] in the case that a Higgs bundle has a non-trivial parabolic structure. Note that because it is based on Fock-Goncharov coordinates, and it is not known whether the approach can be applied to the non-singular case.
@@
In [22], we study the constant $B_{2}$ for the rank $2$ case.
@@ 1.2.2 Limiting configuration
We also note that $\det(h_{\infty})$ induces a $C^{\infty}$-metric of $\det(E)$ on $X$ whose curvature is 00, and hence we may assume that $\det(h_{\infty})=h_{\det(E)}$ by adjusting a positive constant. It is called the limiting configuration. It was introduced by Mazzeo-Swoboda-Weiss-Witt [18] in the rank $2$ case, and by Fredrickson [6] in the higher rank case.
@@ 1.2.3 Convergence
In [18], Mazzeo-Swoboda-Weiss-Witt proved the exponentially rapid convergence of the sequence $h_{t}$ to $h_{\infty}$ as $t\to\infty$ locally on $X^{\circ}$. It is generalized to the higher rank case by Fredrickson in [6] under some mild assumption on the ramification of the projection $\Sigma_{\theta}\to X$ (see [23, Remark 1.2]), and by Szabo and the author in [23] more generally.
@@ 1.3.1 The horizontal direction and the vertical direction
[…]
Here, the first term sits in the degree 00. The hypercohomology group of the complex $\mathop{\rm Def}\nolimits(E,\theta)$ is denoted as $H^{\ast}(X,\mathop{\rm Def}\nolimits(E,\theta))$. It is well known that there exists the following natural isomorphism (see §2.2 for an explanation):
@@
[…]
@@ 1.3.2 $C^{\infty}$-representatives of the horizontal infinitesimal deformations
Let $A^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ denote the complex obtained as the space of global sections of the Dolbeault resolution of $\mathop{\rm Def}\nolimits(E,\theta)$. There exists the line bundle $L$ on $\Sigma_{\theta}$ with an isomorphism $(\pi_{\ast}(L),\pi_{\ast}(\eta))\simeq(E,\theta)$. For any $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$, the multiplication of $\nu$ on $L$ induces $F_{\nu}\in H^{0}(X,\mathop{\rm End}\nolimits(E)\otimes K_{X}(D))$. We would like to modify $F_{\nu}$ to a $1$-cocycle $\tau$ in $A^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ such that the cohomology class $[\tau]$ equals $\iota^{{\tt h}}(\nu)$. We remark that any $1$-cocycle $\tau$ in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ naturally induces a holomorphic section $\rho(\tau)$ of $\pi_{\ast}(K_{\Sigma_{\theta}})$ such that $\Upsilon_{1}([\tau])=\rho(\tau)$. (See §2.2.6.)
@@
A $1$-cocycle $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ satisfies $[\tau]=\iota^{{\tt h}}(\nu)$ if the following holds.
  • There exist a relatively compact open neighbourhood $X_{P}^{\prime}$ of $P$ in $X_{P}$ $(P\in D(\theta))$ such that $\tau_{|X\setminus\bigcup X_{P}^{\prime}}=F_{\nu|X\setminus\bigcup X_{P}^{\prime}}$.
@@ 1.3.3 Harmonic representatives for horizontal infinitesimal deformations
Let $h_{t}$ be the harmonic metric of $(E,t\theta)$ such that $\det(h_{t})=h_{\det(E)}$. Let $g_{X}$ be a conformal metric of $X$. Let ¯$\overline{\partial}_{E}$ denote the Cauchy operator of $E$. The induced Cauchy operator of $\mathop{\rm End}\nolimits(E)$ is also denoted by ¯$\overline{\partial}_{E}$. Let $\iota^{{\tt h}}_{t}:H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})\to H^{1}(X,\mathop{\rm Def}\nolimits(E,t\theta))$ denote the splitting morphism for $(E,t\theta)$, where we identify $\Sigma_{t\theta}$ and $\Sigma_{\theta}$ in a natural way. To study the harmonic representative ${\tt H}(\nu,t)$ of $\iota^{{\tt h}}_{t}(\nu)$, we construct a $1$-cocycle ${\tt H}^{\prime}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that the following holds.
@@
    ${\tt H}^{\prime}(\nu,t)$ satisfies the conditions in Proposition 1.4. In particular, $[{\tt H}^{\prime}(\nu,t)]=\iota^{{\tt h}}_{t}(\nu)$.
@@
    (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t)=O(e^{-\beta t})\|\nu\|$ for some $\beta>0$, where $\|\nu\|$ denotes the $L^{2}$-norm of $\nu$ on $\Sigma_{\theta}$.
@@
Let $P\in D(\theta)$. There exists a holomorphic function $\alpha_{P}$ on $\Sigma_{\theta|X_{P}}$ such that $d\alpha_{P}=\nu_{|\Sigma_{\theta|X_{P}}}$ and that $\alpha_{P}(Q)=0$ for any $Q\in\pi^{-1}(P)$. The multiplication of $\alpha_{P}$ on $L_{|\Sigma_{\theta|X_{P}}}$ induces an endomorphism $F_{\alpha_{P}}$ of $E_{P}$ such that $[\theta,F_{\alpha_{P}}]=0$. Let $h_{P,t}$ be the harmonic metric of $(E_{P},t\theta_{P})$ such that $h_{P,t|\partial X_{P}}=h_{\infty|\partial X_{P}}$. Let $\nabla_{h_{P,t}}=\overline{\partial}_{E}+\partial_{E_{P},h_{P,t}}$ denote the Chern connection of $(E,h_{P,t})$. Let $\theta^{\dagger}_{h_{P,t}}$ denote the adjoint of $\theta$ with respect to $h_{P,t}$. Then, it is easy to see that
@@
(¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t).$
Then, ${\tt H}(\nu,t)={\tt H}^{\prime}(\nu,t)-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)$ are the harmonic representatives of $\iota^{{\tt h}}_{t}(\nu)$. By using the estimate for the first eigenvalues of the Laplacian operators (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)$ (see Proposition 2.42) and the regularity for solutions of the Poisson equation (see Proposition 2.36), we obtain $|\gamma(\nu,t)|_{h_{t}}+|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)|_{h_{t},g_{X}}=O(e^{-\epsilon t})$ for some $\epsilon>0$.

### v3

@@ Abstract
The moduli space of stable Higgs bundles of degree 0 is equipped with the hyperkähler metric, called the Hitchin metric. On the locus where the spectral curves are smooth, there is the hyperkähler metric called the semi-flat metric, associated with the algebraic integrable systems with the Hitchin section. We prove the exponentially rapid decay of the difference between the Hitchin metric and the semi-flat metric along the ray $(E,t\theta)$ as $t\to\infty$.
@@ 1.1.1 Hitchin metric
Let $X$ be a compact Riemann surface with a conformal metric $g_{X}$. In this paper, a Higgs bundle on $X$ is a holomorphic vector bundle $E$ with a holomorphic section $\theta$ of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$, where $K_{X}$ denotes the canonical bundle of $X$. The study of the moduli space of Higgs bundles was initiated by Hitchin in [14]. Let $\mathcal{M}_{H}$ denote the moduli space of stable Higgs bundles of degree 0 and rank $n$ on $X$. It is naturally equipped with the integrable almost complex structure $I$. Moreover, $\mathcal{M}_{H}$ is equipped with the naturally defined hyperkähler structure, i.e., the Riemannian metric $g_{H,{\mathbb{R}}}$ and the three integrable almost complex structures $I$, $J$ and $K$ such that (i) the associated $2$-forms $\omega_{T}(\cdot,\cdot)=g_{H,{\mathbb{R}}}(\cdot,T\cdot)$ $(T=I,J,K)$ are closed, (ii) IJK=1IJK=-1. Let $g_{H}$ denote the associated Kähler metric of $(\mathcal{M}_{H},I)$, i.e., $g_{H}=g_{H,{\mathbb{R}}}+\sqrt{-1}\omega_{I}$. The Riemannian metric $g_{H,{\mathbb{R}}}$ and the Kähler metric $g_{H}$ are called the Hitchin metric.
By the theorem of Hitchin and Simpson [14, 26], because $(E,\theta)\in\mathcal{M}_{H}$ is stable of degree 0, it carries a harmonic metric $h$, i.e., $h$ is a Hermitian metric of $E$ satisfying
@@
Here, $R(h)$ denotes the curvature of the Chern connection of $(E,h)$, and $\theta^{\dagger}_{h}$ denotes the adjoint of $\theta$ with respect to $h$. The induced Hermitian metric on $\mathop{\rm End}\nolimits(E)$ is also denoted by $h$. The tangent space $T_{(E,\theta)}\mathcal{M}$ is identified with the space of harmonic $1$-forms of the harmonic bundle $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$. The Hermitian metric $(g_{H})_{|(E,\theta)}$ is identified with the $L^{2}$-metric on the space of harmonic $1$-forms. (For example, see §4.2.3.) Namely, for harmonic $1$-forms $\tau$ of $(\mathop{\rm End}\nolimits(E),\mathop{\rm ad}\nolimits\theta,h)$, let $\tau]\in T_{(E,\theta)}\mathcal{M}_{H}$ denote the corresponding element, and then we have
@@ 1.1.2 Semi-flat metric
We set $\mathcal{A}_{H}=\bigoplus_{j=1}^{n}H^{0}(X,K_{X}^{j})$. Let $\pi_{0}:T^{\ast}X\to X$ denote the projection. Let $\eta$ denote the tautological section of $\pi_{0}^{\ast}(K_{X})$. For any $a=(a_{j})\in\mathcal{A}_{H}$, we obtain the section $\xi_{a}=\eta^{n}+\sum_{j=1}^{n}(-1)^{j}a_{j}\eta^{n-j}$ of $\pi_{0}^{\ast}(K_{X}^{\otimes n})$. Let $\Sigma_{a}$ denote the 0-scheme of the section $\xi_{a}$. Such a curve is called a spectral curve. Let $\mathcal{A}_{H}^{\prime}\subset\mathcal{A}_{H}$ denote the Zariski open subset of $a\in\mathcal{A}_{H}$ such that $\Sigma_{a}$ are smooth. It is known that $\mathcal{A}_{H}^{\prime}$ is non-empty.
@@
It is independent of the choice of a Hitchin section. Because the fiber $\Phi_{H}^{-1}(\Phi_{H}(E,\theta))$ is identified with a connected component of the Picard variety of $\Sigma_{\theta}$, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt v}}$ is naturally identified with $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})$. The moduli space $\mathcal{M}_{H}^{\prime}$ is equipped with a complex symplectic structure. Because $\Phi_{H}^{-1}(\Phi_{H}(E,\theta))$ is Lagrangian, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt v}}$ and $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt h}}$ are mutually dual. Hence, $(T_{(E,\theta)}\mathcal{M}_{H}^{\prime})^{{\tt h}}$ is naturally identified with $H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$. The space $H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\oplus H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ is identified with the space of harmonic $1$-forms on $\Sigma_{\theta}$. The Hermitian metric $(g_{\mathop{\rm sf}\nolimits})_{(E,\theta)}$ is identified with the $L^{2}$-metric on the space of harmonic $1$-forms on $\Sigma_{\theta}$. (See Proposition 4.6.) Namely, for harmonic $1$-forms $\tau$ on $\Sigma_{\theta}$, let $\tau]\in H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})\oplus H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$ denote the corresponding element, and then we have
@@ 1.1.3 A weak version of a conjecture of Gaiotto-Moore-Neitzke
A previous work was done by Fredrickson in [5] on the basis of her another work [6], the works of Mazzeo-Swoboda-Weiss-Witt [18, 19] and interesting ideas of Dumas-Neitzke [4]. More precisely, for each $(E,\theta)\in\mathcal{M}_{H}^{\prime}$, let $\mathop{\rm Crit}\nolimits(\theta)\subset\Sigma_{\theta}$ denote the set of the critical points of the projection $\pi:\Sigma_{\theta}\to X$. For each $Q\in\mathop{\rm Crit}\nolimits(\theta)$, let $r_{Q}$ denote the ramification index of $\pi$ at $Q$. We obtain the Zariski open subset $\mathcal{M}_{H}^{\prime\prime}$ of $(E,\theta)\in\mathcal{M}_{H}^{\prime}$ such that $r_{Q}\leq 2$ for any $Q\in\mathop{\rm Crit}\nolimits(\theta)$. We have $\mathcal{M}^{\prime}_{H}=\mathcal{M}_{H}^{\prime\prime}$ if $n=2$ but $\mathcal{M}_{H}^{\prime}\neq\mathcal{M}_{H}^{\prime\prime}$ if $n\geq 3$, in general. The conjectural estimate (1) has been studied for any $(E,\theta)\in\mathcal{M}_{H}^{\prime\prime}$ in [5]. We should also note that a similar estimate has been studied by Fredrickson-Mazzeo-Swoboda-Weiss [7] in the case of parabolic Higgs bundles of rank $2$. Chen-Li [2] studied the rank $2$ but more singular case by using the methods in [5, 7] and this paper. In [15], Holdt studied the original approach of [9, 10] in the case that a Higgs bundle has a non-trivial parabolic structure. Note that because it is based on Fock-Goncharov coordinates, it is not known whether the approach can be applied to the non-singular case.
@@
In [22], we study the constant $B_{2}$ more closely for the rank $2$ case. In [23], we study a generalization to the non-zero degree case.
@@ 1.2.2 Limiting configuration
We also note that $\det(h_{\infty})$ induces a $C^{\infty}$-metric of $\det(E)$ on $X$ whose curvature is 0, and hence we may assume that $\det(h_{\infty})=h_{\det(E)}$ by adjusting a positive constant. It is called the limiting configuration. It was introduced by Mazzeo-Swoboda-Weiss-Witt [18] in the rank $2$ case, and by Fredrickson [6] in the higher rank case.
@@ 1.2.3 Convergence
In [18], Mazzeo-Swoboda-Weiss-Witt proved the exponentially rapid convergence of the sequence $h_{t}$ to $h_{\infty}$ as $t\to\infty$ locally on $X^{\circ}$. It is generalized to the higher rank case by Fredrickson in [6] under some mild assumption on the ramification of the projection $\Sigma_{\theta}\to X$ (see [24, Remark 1.2]), and by Szabo and the author in [24] more generally.
@@ 1.3.1 The horizontal direction and the vertical direction
$\begin{CD}\mathop{\rm End}\nolimits(E)@>{\mathop{\rm ad}\nolimits\theta}>{}>\mathop{\rm End}\nolimits(E)\otimes K_{X}.\end{CD}$
Here, the first term sits in the degree 0. The hypercohomology group of the complex $\mathop{\rm Def}\nolimits(E,\theta)$ is denoted as $H^{\ast}(X,\mathop{\rm Def}\nolimits(E,\theta))$. It is well known that there exists the following natural isomorphism (see §2.2 for an explanation):
@@
$\begin{CD}0@>{}>{}>H^{1}(\Sigma_{\theta},\mathcal{O}_{\Sigma_{\theta}})@>{\iota^{{\tt v}}}>{}>H^{1}(X,\mathop{\rm Def}\nolimits(E,\theta))@>{\Upsilon_{1}}>{}>H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})@>{}>{}>0.\end{CD}$
@@ 1.3.2 $C^{\infty}$-representatives of the horizontal infinitesimal deformations
Let $A^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ denote the complex obtained as the space of global sections of the Dolbeault resolution of $\mathop{\rm Def}\nolimits(E,\theta)$. There exists the line bundle $L$ on $\Sigma_{\theta}$ with an isomorphism $(\pi_{\ast}(L),\pi_{\ast}(\eta))\simeq(E,\theta)$. For any $\nu\in H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})$, the multiplication of $\nu$ on $L$ induces $F_{\nu}\in H^{0}(X,\mathop{\rm End}\nolimits(E)\otimes K_{X}(D))$. We would like to modify $F_{\nu}$ to a $1$-cocycle $\tau$ in $A^{\bullet}(\mathop{\rm Def}\nolimits(E,\theta))$ such that the cohomology class $\tau]$ equals $\iota^{{\tt h}}(\nu)$. We remark that any $1$-cocycle $\tau$ in $A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ naturally induces a holomorphic section $\rho(\tau)$ of $\pi_{\ast}(K_{\Sigma_{\theta}})$ such that $\Upsilon_{1}([\tau])=\rho(\tau)$. (See §2.2.6.)
@@
A $1$-cocycle $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta))$ satisfies $\tau]=\iota^{{\tt h}}(\nu)$ if the following holds.
  • For each $P\in D(\theta)$, there exists a relatively compact open neighbourhood $X_{P}^{\prime}$ of $P$ in $X_{P}$ such that $\tau_{|X\setminus\bigcup X_{P}^{\prime}}=F_{\nu|X\setminus\bigcup X_{P}^{\prime}}$.
@@ 1.3.3 Harmonic representatives for horizontal infinitesimal deformations
Let $h_{t}$ be the harmonic metric of $(E,t\theta)$ such that $\det(h_{t})=h_{\det(E)}$. Let $g_{X}$ be a conformal metric of $X$. Let $\overline{\partial}_{E}$ denote the Cauchy operator of $E$. The induced Cauchy operator of $\mathop{\rm End}\nolimits(E)$ is also denoted by $\overline{\partial}_{E}$. Let $\iota^{{\tt h}}_{t}:H^{0}(\Sigma_{\theta},K_{\Sigma_{\theta}})\to H^{1}(X,\mathop{\rm Def}\nolimits(E,t\theta))$ denote the splitting morphism for $(E,t\theta)$, where we identify $\Sigma_{t\theta}$ and $\Sigma_{\theta}$ in a natural way. To study the harmonic representative ${\tt H}(\nu,t)$ of $\iota^{{\tt h}}_{t}(\nu)$, we construct a $1$-cocycle ${\tt H}^{\prime}(\nu,t)\in A^{1}(\mathop{\rm Def}\nolimits(E,t\theta))$ such that the following holds.
@@
    ${\tt H}^{\prime}(\nu,t)$ satisfies the conditions in Proposition 1.4. In particular, ${\tt H}^{\prime}(\nu,t)]=\iota^{{\tt h}}_{t}(\nu)$.
@@
    $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t)=O(e^{-\beta t})\|\nu\|$ for some $\beta>0$, where $\|\nu\|$ denotes the $L^{2}$-norm of $\nu$ on $\Sigma_{\theta}$.
@@
Let $P\in D(\theta)$. There exists a holomorphic function $\alpha_{P}$ on $\Sigma_{\theta|X_{P}}$ such that $d\alpha_{P}=\nu_{|\Sigma_{\theta|X_{P}}}$ and that $\alpha_{P}(Q)=0$ for any $Q\in\pi^{-1}(P)$. The multiplication of $\alpha_{P}$ on $L_{|\Sigma_{\theta|X_{P}}}$ induces an endomorphism $F_{\alpha_{P}}$ of $E_{P}$ such that $\theta,F_{\alpha_{P}}]=0$. Let $h_{P,t}$ be the harmonic metric of $(E_{P},t\theta_{P})$ such that $h_{P,t|\partial X_{P}}=h_{\infty|\partial X_{P}}$. Let $\nabla_{h_{P,t}}=\overline{\partial}_{E}+\partial_{E_{P},h_{P,t}}$ denote the Chern connection of $(E,h_{P,t})$. Let $\theta^{\dagger}_{h_{P,t}}$ denote the adjoint of $\theta$ with respect to $h_{P,t}$. Then, it is easy to see that
@@
$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)=(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}{\tt H}^{\prime}(\nu,t).$
Then, ${\tt H}(\nu,t)={\tt H}^{\prime}(\nu,t)-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)$ are the harmonic representatives of $\iota^{{\tt h}}_{t}(\nu)$. By using the estimate for the first eigenvalues of the Laplacian operators $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)^{\ast}_{h_{t},g_{X}}\circ(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)$ (see Proposition 2.41) and the regularity for solutions of the Poisson equation (see Proposition 2.35), we obtain $|\gamma(\nu,t)|_{h_{t}}+|(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits t\theta)\gamma(\nu,t)|_{h_{t},g_{X}}=O(e^{-\epsilon t})$ for some $\epsilon>0$.